OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8

OpenStudy (hoblos):

replace x by 4 you get g(4) = 4^2 + 3 = ?

OpenStudy (anonymous):

19?

OpenStudy (whpalmer4):

Yes, 19 is correct. \[g(4) = 4^2+3 = 4*4 + 3 = 16+3 = 19\]